# One Perturbation is Enough: On Generating Universal Adversarial Perturbations against Vision-Language Pre-training Models

### v2

@@ Abstract
Vision-Language Pre-training (VLP) models have exhibited unprecedented capability in many applications by taking full advantage of the multimodal alignment. However, previous studies have shown they are vulnerable to maliciously crafted adversarial samples. Despite recent success, these methods are generally instance-specific and require generating perturbations for each input sample. In this paper, we reveal that VLP models are also vulnerable to the instance-agnostic universal adversarial perturbation (UAP). Specifically, we design a novel Contrastive-training Perturbation Generator with Cross-modal conditions (C-PGC) to achieve the attack. In light that the pivotal multimodal alignment is achieved through the advanced contrastive learning technique, we devise to turn this powerful weapon against themselves, i.e., employ a malicious version of contrastive learning to train the C-PGC based on our carefully crafted positive and negative image-text pairs for essentially destroying the alignment relationship learned by VLP models. Besides, C-PGC fully utilizes the characteristics of Vision-and-Language (V+L) scenarios by incorporating both unimodal and cross-modal information as effective guidance. Extensive experiments show that C-PGC successfully forces adversarial samples to move away from their original area in the VLP model’s feature space, thus essentially enhancing attacks across various victim models and V+L tasks. The GitHub repository is available at https://github.com/ffhibnese/CPGC_VLP_Universal_Attacks.
@@ 1 Introduction
Vision-Language Pre-training (VLP) models, including ALBEF (Li et al., 2021), TCL (Yang et al., 2022), and BLIP (Li et al., 2022), have recently demonstrated remarkable efficacy in a wide range of Vision-and-Language (V+L) tasks. By self-supervised pre-training on large-scale image-text pairs, VLP models efficiently align cross-modal features and capture rich information from the aligned multimodal embeddings, thereby providing expressive representations for various applications.
@@
Adversarial attacks (Carlini & Wagner, 2017), which aim to deceive models during inference time, have attracted extensive attention due to their significant threat to security-critical scenarios (Eykholt et al., 2018). Recent studies have shown that VLP models are also vulnerable to adversarial samples. The pioneering work Co-Attack (Zhang et al., 2022) proposes the first multimodal attack that simultaneously perturbs both image and text modalities and displays excellent performance. However, Co-Attack only considers relatively easier white-box attacks where victim models are completely accessible. To handle more practical black-box settings, subsequent studies propose various transferable adversarial samples generated on an available surrogate model to fool other inaccessible models. Specifically, SGA (Lu et al., 2023) significantly improves the adversarial transferability through the set-level cross-modal guidance obtained from data augmentations. Subsequently, TMM (Wang et al., 2024) proposes to jointly destroy the modality-consistency features within the clean image-text pairs and include more modality-discrepancy features in the perturbations to further enhance transferability. While existing methods have achieved great success, they are all instance-specific and need to generate a perturbation for each input pair, which results in substantial computational overhead. Meanwhile, universal adversarial attacks, as an efficient instance-agnostic approach that uses only one Universal Adversarial Perturbation (UAP) to conduct attacks, have not been fully investigated for VLP models. This naturally leads to a question, is it possible to design a universal adversarial perturbation that can effectively deceive VLP models across various image-text pairs?
@@
Motivation. To answer this question, we make an intuitive attempt to transplant existing renowned approaches UAP (Moosavi-Dezfooli et al., 2017) and GAP (Poursaeed et al., 2018) to launch attacks on several VLP models by maximizing the distance between the embeddings of the adversarial image and its matched texts (Lu et al., 2023). Unfortunately, Figure 1 demonstrates that these methods yield unsatisfactory attack success rates (ASR), especially for black-box attacks. Empirically, this failure stems from their narrow focus on the image modality, disregarding the other modality and the multimodal information that plays a pivotal role in VLP models. To overcome this challenge, we revisit the VLP models’ basic training paradigm and emphasize that regardless of the downstream V+L tasks, the outstanding performance achieved by these models is heavily reliant on the successful vision-language feature alignment, which draws the embedding of matched image-text pairs closer while distancing those of non-matched pairs. In light of this consideration, we argue that the key core of an effective universal adversarial attack is to obtain a UAP that can fundamentally destroy this learned alignment relationship to mislead VLP models into making incorrect decisions. In addition, Fig. 1 also shows that the generator-based approach GAP consistently achieves superior ASR than UAP, confirming the validity of the generative attack paradigm (Hashemi et al., 2020).
@@
Based on these insights, we propose a generative UAP framework that learns a Contrastive-training Perturbation Generator with Cross-modal conditions (C-PGC) to launch attacks on VLP models (see Fig. 2). To essentially destroy the multimodal alignment, we devise to utilize VLP models’ most powerful weapons to attack against themselves, i.e., utilize the contrastive learning mechanism to train the generator using our maliciously constructed image-text pairs that completely violate the correct image-text matching relationship, to produce perturbation that pushes the embedding of matched pairs apart while pulling those of non-matched ones together. Inspired by the multimodal characteristics of V+L scenarios, we modify the generator’s architecture to incorporate cross-modal knowledge through the advanced cross-attention mechanism for better guidance. In addition, we also consider the intra-modal influence and introduce a unimodal distance loss to further enhance the attacks. Since previous studies (Zhang et al., 2022; Lu et al., 2023) achieve impressive improvements via multimodal perturbation, we are motivated to generate UAP for both images and texts to utilize the synergy between different modalities. Our contributions can be summarized as follows:
@@
We design a perturbation generator conditioned with cross-modal knowledge to produce effective UAP added to both image and text modalities for universal adversarial attacks on VLP models.
We propose a novel training framework that incorporates both unimodal and multimodal guidance to contrastively train the generator using our meticulously constructed positive and negative pairs.
@@
Extensive experiments on 6 various VLP models across different V+L tasks reveal that our method achieves outstanding white-box performance and black-box transferability in different scenarios.
@@ 2.1 Vision-Language Pre-training Models
VLP models are pre-trained on massive image-text pairs to learn the semantic correlations across modalities and serve diverse multimodal user demands (Chen et al., 2023; Du et al., 2022). We next illustrate the basis of VLP models from multiple perspectives.
Architectures. Based on the ways of multimodal fusion, the architectures of VLP models can be classified into two types: single-stream and dual-stream architectures. Single-stream architectures (Li et al., 2019; Chen et al., 2020) directly concatenate the text and image features, and calculate the attention in the same Transformer block for multimodal fusion. On the contrary, dual-stream architectures (Radford et al., 2021; Li et al., 2022) separately feed the text and image features to different Transformer blocks and leverage the cross-attention mechanism for multimodal fusion.
Pre-training Objectives. The pre-training objectives for VLP models mainly include masked features completion, multimodal features matching, and specific downstream objectives. Masked features completion (Chen et al., 2020) encourages VLP models to predict the deliberately masked tokens using the remaining unmasked tokens during pre-training. Multimodal features matching (Li et al., 2021) pre-trains VLP models by learning to precisely predict whether the given image-text pairs are matched. Specific downstream objectives (Anderson et al., 2018) directly utilize the training objectives of downstream tasks (e.g., visual question answering) for pre-training VLP models.
Downstream Tasks. In this paper, we mainly consider the following multimodal downstream tasks: (1) Image-text retrieval (ITR) (Wang et al., 2016): finding the most matched image for the given text and vice versa, including image-to-text retrieval (TR) and text-to-image retrieval (IR). (2) Image caption (IC) (Bai & An, 2018): generating the most suitable descriptions for the given image. (3) Visual grounding (VG) (Hong et al., 2019): locating specific regions in the image that correspond with the given textual descriptions. (4) Visual entailment (VE) (Xie et al., 2019): analyzing the input image and text and predicting whether their relationship is entailment, neutral, or contradiction.
@@ 2.2 Adversarial Attacks
Instance-specific Attacks on VLP Models. The adversarial robustness of VLP Models has already become a research focus. Early works (Kim & Ghosh, 2019; Yang et al., 2021) impose perturbations only on single modality and lack cross-modal interactions when attacking multimodal models. To address this issue, Co-Attack (Zhang et al., 2022) conducts the first multimodal white-box attacks on VLP models. On the basis of Co-Attack, Lu et al. (2023) extend the attacks to more rigorous black-box settings and propose SGA, which utilizes set-level alignment-preserving argumentations with carefully designed cross-modal guidance. However, Wang et al. (2024) points out that SGA fails to fully exploit modality correlation, and proposes TMM to better leverage cross-modal interactions by tailoring both the modality-consistency and modality-discrepancy features. Nonetheless, these methods are all instance-specific and need to craft perturbations for each input pair.
Universal Adversarial Examples. Universal adversarial attacks (Moosavi-Dezfooli et al., 2017; Mopuri et al., 2018) aim to deceive the victim model by exerting a uniform adversarial modification to all the benign samples. These attacks save the redundant procedures of redesigning perturbations for each input sample and are consequently more efficient than traditional attack strategies. Generally, universal adversarial attacks can be categorized into optimization-based methods (Moosavi-Dezfooli et al., 2017; Wang et al., 2023; Liu et al., 2023) and generation-based methods (Hayes & Danezis, 2018; Gao et al., 2024; Anil et al., 2024). Benefiting from the powerful modeling abilities of generative models, generation-based methods are more versatile and can produce more natural samples than optimization-based ones. In this paper, we explore universal adversarial attacks on VLP models and manage to generate UAP with excellent attack effects and high transferability.
@@ 3 Universal Multimodal Attacks
In this section, we first present the problem statement of universal adversarial attacks on VLP models. Next, we introduce the overview of our framework. Finally, we illustrate the detailed design of each proposed technique and summarize the training objective and paradigm of C-PGC.
@@ 3.1 Problem Statement
We define an input image-text pair as and denote and as the image and text embedding encoded by the image encoder and text encoder of the targeted VLP model . Let be an available dataset consisting of image-text pairs collected by a malicious adversary. The attack objective is to utilize to train a generator that is capable of producing a powerful pair of universal image-text perturbations that can affect the vast majority of test dataset to fool models into making incorrect decisions. Formally, the attack goal can be formulated as:
where denotes the operation that uses the output V+L features to obtain the final predictions, indicates the text perturbation strategy (Zhang et al., 2022; Lu et al., 2023) that replaces certain important tokens of the original sentence with crafted adversarial words, and is the correct prediction of the considered V+L task. To ensure the perturbation’s imperceptibility, we constrain the pixel-level image perturbation with norm of a given budget . The textual perturbation is token-level and the stealthiness is accordingly constrained by the number of modified words . Since altering words in a natural sentence can be easily noticed or detected, we apply a rigorous restriction that permits only a single word to be substituted (). On the premise of imperceptibility, the attacker attempts to generalize the crafted UAP to a wider range of test data and victim models.
@@ 3.2 Overview of the Proposed Framework
As depicted in Fig. 3, we adopt the multimodal perturbation strategy and generate perturbations on both image and text modalities for enhanced attacks. Given the similarity between the workflows for image and text, we then take the image attacks as an example to illustrate the proposed framework.
Firstly, a fixed noise is randomly initialized and subsequently fed into the conditional generator. For each image and its descriptions , the generator then translates the input noise into the adversarial perturbation that is of the same size as the image . During the generation, the network additionally benefits from cross-modal information by integrating the textual embedding, i.e., . Next, the generated adversarial noise is injected into the clean image to obtain the adversarial image via . To better guide the training process, we design two efficient unimodal and multimodal losses as our optimization objectives. Unimodal loss is straightforward and aims to push the adversarial images away from the clean images in the latent embedding space, while multimodal loss is based on contrastive learning and uses our manually constructed positive and negative samples to effectively destroy the image-text matching relationship obtained from feature alignment. Once we finish training the C-PGC using the proposed losses, the input fixed noise is transformed into a UAP that is of great generalization and transferability.
@@ 3.3 Detailed Design of C-PGC
Perturbation Generator with Cross-modal conditions. Previous generative universal attacks (Gao et al., 2024; Anil et al., 2024) have shown excellent efficacy in fooling the discriminative models. Nevertheless, since existing generative attacks are limited to a single modality, directly utilizing the off-the-shelf generators might fail to leverage the multimodal interactions in these special V+L scenarios. To address this limitation, we additionally introduce cross-modal embeddings as auxiliary information to further facilitate the process of perturbation generation. Specifically, we modify the existing generator’s architecture by adding several cross-attention modules that have been proven effective in tasks with variable input modalities. The obtained textual embeddings encoded by are then incorporated into our generator through:
@@
where is the flattened intermediate features within the generator, , , are optimized parameters in the attention modules.
@@
Multimodal Contrastive Loss. The preceding analysis regarding the failures of existing UAP attacks encourages us to design a loss function that can guide the generated UAP to break the learned multimodal feature alignment. Motivated by the fact that contrastive learning underpins the cross-modal alignment, we advocate leveraging this mechanism to attack VLP models themselves by contrastively training our C-PGC to essentially disrupt the benign alignment relationship. Concretely, we adopt the widely recognized InfoNCE (He et al., 2020) as our basic contrastive loss.
@@
To establish the contrastive paradigm, we first define the adversarial image as the anchor sample. Besides, it is also necessary to construct an appropriate set of positive and negative samples. Based on the fundamental objective of our attacks, it is natural that the originally matched text descriptions set can be employed as negative samples to increase the discrepancy of matched image-text in the feature space of VLP models. To further push the adversarial image away from its correct text descriptions, we propose utilizing multiple texts whose embeddings differ significantly from that of the original clean images as the positive samples . Specifically, we randomly sample a batch of text sets from and select the text set with the farthest feature distances from the current image as positive samples, i.e., . Moreover, we utilize data augmentations that resize the clean image into different scales and apply random Gaussian noise to acquire a more diverse image set for set-level guidance Lu et al. (2023). With the augmented image set and these well-constructed positive and negative samples, the multimodal contrastive loss can be formulated as:
@@
where is the universal image perturbation, denotes the temperature parameter and , where represents the cosine similarity measurement.
@@
Unimodal Distance Loss. Apart from the multimodal guidance, we also consider the unimodal influence by directly pushing adversarial images away from their initial visual semantic area to further enhance attack effects. Similarly, to acquire set-level guidance, the input image is initially resized to different scales and then added with Gaussian noise to generate the augmented image set . Then, we craft the adversarial image through and process with the same augmentation operation to obtain the adversarial image set . Finally, we minimize the negative Euclidean distance between the embeddings of adversarial images and clean images to optimize the UAP generator. Formally, the loss is formulated as:
@@
Taking advantage of the unimodal set-level guidance, ensures an effective optimization direction during the generator training and further improves the attack effectiveness of our UAP.
@@
Training Objective. With the two well-designed loss terms and , the overall optimization objective of our conditional generator concerning image attacks can be formulated as:
@@
where is the pre-defined hyperparameter to balance the contributions of and . By training the network using the proposed loss function based on the data distribution of the multimodal training dataset , the generator is optimized to produce UAP that can push the features of mismatched image-text pairs together while pulling the embeddings of the matched ones apart, thereby learning a UAP with excellent generalization ability and adversarial transferability.
@@
Text Modality Attacks. In textual attacks, the UAP generator’s architecture and training loss are completely symmetrical with those of image attacks. Correspondingly, embeddings of the matched image are used as the cross-modal conditions for the generator. Given an adversarial text as the anchor sample, we use the set scaled from the originally matched image as negative samples while the augmented from the farthest image within the randomly sampled image set as positive samples to formulate the loss. is consequently calculated as the negative Euclidean distance between the embeddings of and the clean input .

### v3

@@ Abstract
Vision-Language Pre-training (VLP) models have exhibited unprecedented capability in many applications by taking full advantage of the learned multimodal alignment. However, previous studies have shown they are vulnerable to maliciously crafted adversarial samples. Despite recent success, these attacks are generally instance-specific and require generating perturbations for each input sample. In this paper, we reveal that VLP models are also susceptible to the instance-agnostic universal adversarial perturbation (UAP). Specifically, we design a novel Contrastive-training Perturbation Generator with Cross-modal conditions (C-PGC). In light that the pivotal multimodal alignment in VLP models is achieved via contrastive learning, we devise to turn this powerful weapon against VLP models themselves. I.e., we employ a malicious version of contrastive learning to train the proposed generator using our carefully crafted positive and negative image-text pairs. Once training is complete, the generator is able to produce universal perturbations that can essentially destroy the established alignment relationship in VLP models. Besides, C-PGC fully utilizes the characteristics of Vision-and-Language (V+L) scenarios by incorporating both unimodal and cross-modal information as effective guidance. Extensive experiments show that C-PGC successfully forces adversarial samples to move away from their original area in the VLP model’s feature space, thus fundamentally enhancing attack performance across various victim models and V+L tasks.
@@ 1 Introduction
Vision-Language Pre-training (VLP) models have recently demonstrated remarkable efficacy in a wide range of Vision-and-Language (V+L) tasks. By self-supervised pre-training on large-scale image-text pairs, VLP models efficiently align cross-modal features and capture rich information from the aligned multimodal embeddings, thereby providing expressive representations for various applications.
@@
Adversarial attacks [7], which aim to deceive models during inference time, have attracted extensive attention due to their significant threat to security-critical scenarios [15]. Recent studies have shown that VLP models are also vulnerable to adversarial samples. The pioneering work Co-Attack [48] proposes the first multimodal attack that simultaneously perturbs both image and text modalities and displays excellent performance. However, Co-Attack only considers relatively easier white-box attacks where victim models are completely accessible. To handle more practical black-box settings, subsequent studies propose various transferable adversarial samples generated on an available surrogate model to fool other inaccessible models. Specifically, SGA [29] significantly improves the adversarial transferability through the set-level cross-modal guidance obtained from data augmentations. Subsequently, TMM [41] proposes to jointly destroy the modality-consistency features within the clean image-text pairs and include more modality-discrepancy features in the perturbations to further enhance transferability. While existing methods have achieved great success, they are all instance-specific and need to generate a perturbation for each input pair, which results in substantial computational overhead. Meanwhile, universal adversarial attacks (illustrated in Fig. 1), as an efficient instance-agnostic approach that uses only one Universal Adversarial Perturbation (UAP) to conduct attacks, have not been fully investigated for VLP models. This naturally leads to a question, is it possible to design a UAP that can deceive VLP models across different image-text pairs?
@@
Motivation. To this end, we make an intuitive attempt to transplant renowned approaches UAP [31] and GAP [37] to attack several VLP models by maximizing the distance between the embeddings of the adversarial image and its matched texts. Unfortunately, Fig. 2 demonstrates that these methods yield unsatisfactory attack success rates (ASR), especially for black-box attacks. Empirically, this failure stems from their narrow focus on the image modal, disregarding the other modality and the multimodal information that plays a pivotal role in VLP models. To overcome this challenge, we revisit VLP models’ basic training paradigm and emphasize that regardless of the downstream V+L tasks, their achieved outstanding performance is heavily reliant on the well-established multimodal alignment, which draws the embeddings of matched image-text pairs closer while distancing those of non-matched pairs. In light of this consideration, we argue that the core of an effective universal adversarial attack is to obtain a UAP that can fundamentally destroy this learned alignment relationship to mislead VLP models into making incorrect decisions. Besides, Fig. 2 shows that the generator-based GAP consistently outperforms UAP methods, due to the excellent distributional modeling capability of generators. This suggests the superiority of the generative paradigm, which is also corroborated by numerous studies [17, 16].
@@
Based on these insights, we propose a novel generative framework that learns a Contrastive-training Perturbation Generator with Cross-modal conditions (C-PGC) to launch universal attacks on VLP models. To essentially destroy the multimodal alignment, we devise to utilize VLP models’ most powerful weapons to attack against themselves. I.e., utilize the contrastive learning mechanism to train the generator based on maliciously constructed image-text pairs that completely violate the correct V+L matching relationship, to produce perturbation that pushes the embeddings of matched pairs apart while pulling those of non-matched ones together. Moreover, most previous studies [48, 29, 49] simply exploit the crucial cross-modal information by maximizing the feature distance between samples of different modals to optimize perturbations, without deeper exploration for attack enhancement. In contrast, we fully harness V+L characteristics by refining the generator architecture to incorporate cross-modal knowledge through cross-attention mechanisms for better guidance. Besides, we also consider the intra-modal influence and introduce a unimodal distance loss to further improve attack effectiveness. We highlight that the proposed framework is seamlessly compatible with text perturbation generation, achieving a truly multimodal universal attack that benefits from the synergy between V+L modalities. Our contributions are as follows:
@@
We propose C-PGC, a novel perturbation generator conditioned on cross-modal knowledge, to produce both image and text UAPs for powerful attacks on VLP models.
We design a malicious contrastive learning paradigm that incorporates both unimodal and multimodal guidance to train the generator to produce UAP that can essentially disrupt the multimodal alignment in VLP models.
@@
Extensive experiments on 6 various VLP models across diverse V+L tasks reveal that our method achieves outstanding attack performance in different scenarios.
@@ 2.1 Vision-Language Pre-training Models
VLP models are pre-trained on massive image-text pairs to learn the semantic correlations across modalities and serve diverse multimodal user demands [8, 13]. We next illustrate the basis of VLP models from multiple perspectives.
Architectures. Based on the ways of multimodal fusion, the architectures of VLP models can be classified into two types: single-stream and dual-stream architectures. Single-stream architectures [25, 9] directly concatenate the text and image features, and calculate the attention in the same Transformer block for multimodal fusion. On the contrary, dual-stream architectures [38, 24] separately feed the text and image features to different Transformer blocks and leverage the cross-attention mechanism for multimodal fusion. Generally, single-stream architectures are more parameter-efficient than dual-stream architectures since they adopt the same set of parameters in a Transformer block for the text and image modalities.
Pre-training Objectives. The pre-training objectives for VLP models mainly include masked feature completion, multimodal feature matching, and specific downstream objectives. Masked feature completion [9] encourages VLP models to predict the deliberately masked tokens based on the remaining unmasked tokens. Multimodal feature matching [23] pre-trains VLP models to precisely predict whether the given image-text pairs are matched. Specific downstream objectives [2] directly utilize the training objectives of downstream tasks (e.g., visual question answering) for pre-training VLP models.
Downstream Tasks. In this paper, we mainly consider the following multimodal downstream tasks: (1) Image-text retrieval (ITR) [42]: finding the most matched image for the given text and vice versa, including image-to-text retrieval (TR) and text-to-image retrieval (IR). (2) Image caption (IC) [4]: generating the most suitable descriptions for the given image. (3) Visual grounding (VG) [20]: locating specific regions in the image that correspond with the given textual descriptions. (4) Visual entailment (VE) [43]: analyzing the input image and text and predicting whether their relationship is entailment, neutral, or contradiction.
@@ 2.2 Adversarial Attacks
Instance-specific Attacks on VLP Models. The adversarial robustness of VLP Models has already become a research focus. Early works [22, 45] impose perturbations only on single modality and lack cross-modal interactions when attacking multimodal models. To address this issue, Co-Attack [48] conducts the first multimodal white-box attacks on VLP models. On the basis of Co-Attack, [29] extends the attacks to more rigorous black-box settings and proposes SGA, which utilizes set-level alignment-preserving argumentations with carefully designed cross-modal guidance. However, [41] points out that SGA fails to fully exploit modality correlation, and proposes TMM to better leverage cross-modal interactions by tailoring both the modality-consistency and modality-discrepancy features. Nonetheless, these methods are all instance-specific and need to craft perturbations for each input pair.
Universal Adversarial Examples. Universal adversarial attacks [31, 32, 51] aim to deceive the victim model by exerting a uniform adversarial perturbation to all samples. These attacks save the redundant procedures of redesigning perturbations for each input sample and are hence more efficient than instance-specific methods. Generally, universal adversarial attacks can be categorized into optimization-based [31, 40, 28] and generation-based [18, 17, 3, 51] methods. Benefiting from the powerful modeling abilities of generative models, generation-based methods are more versatile and can produce more natural samples than optimization-based ones. A concurrent work ETU [49] also investigates UAP on VLP models and proposes a data augmentation named ScMix. However, ETU adopts a non-generative approach that narrowly focuses on image UAP, failing to constitute a truly multimodal attack for VLP models. Moreover, ETU demonstrates insufficient attack effects, especially for black-box transferability. In contrast, we propose a generative multimodal attack framework based on a malicious variant of contrastive learning, which yields UAP with strong attack effects and high transferability.
@@ 3 Universal Multimodal Attacks
In this section, we first present the problem statement of universal adversarial attacks on VLP models. Next, we introduce the overview of our framework. Finally, we illustrate the detailed design of the proposed C-PGC.
@@ 3.1 Problem Statement
We define an input image-text pair as and denote and as the image and text embeddings encoded by the image encoder and text encoder of the targeted VLP model . Let be an available dataset consisting of image-text pairs collected by a malicious adversary. The attack objective is to utilize to train a generator for producing a powerful pair of universal image-text perturbations that can affect the vast majority of test dataset to fool models into making incorrect decisions. Formally, the attack goal can be formulated as:
where denotes the operation that uses the output V+L features to obtain the final predictions, indicates the text perturbation strategy [48, 29] that replaces certain important tokens of the original sentence with crafted adversarial words, and is the correct prediction of the considered V+L task. To ensure the perturbation’s imperceptibility, we constrain the pixel-level image perturbation with norm of a given budget . Following previous studies [48, 29, 41], the textual perturbation is token-level and the stealthiness is accordingly constrained by the number of modified words . To ensure the stealthiness of text perturbation, we apply a rigorous restriction that permits only a single word to be substituted (). On the premise of imperceptibility, the attacker attempts to generalize the crafted UAP to a wider range of test data and victim models.
@@ 3.2 Overview of the Proposed Framework
The overview of C-PGC is depicted in Fig. 3. We adopt a multimodal perturbation strategy and generate perturbations for both image and text modalities. Given the high similarity between the workflows of image and text, we then take image attacks as an example for illustration.
Firstly, a fixed noise is randomly initialized and subsequently fed into the conditional generator. For each image and its descriptions , the generator translates the input noise into the adversarial perturbation that is of the same size as . During generation, the network additionally benefits from cross-modal information by integrating the embedding of text descriptions corresponding to the current input image , i.e., . Next, the generated adversarial noise is injected into the clean image to obtain the adversarial image via . To better guide the training process, we design two effective unimodal and multimodal losses as our optimization objectives. Unimodal loss is straightforward and aims to push the adversarial images away from the clean images in the latent embedding space, while multimodal loss is based on contrastive learning using our manually constructed positive and negative samples to strongly destroy the image-text matching relationship achieved by feature alignment. Once we finish training C-PGC using the proposed loss function, the input fixed noise is transformed into a UAP with great generalization and transferability.
@@ 3.3 Detailed Design of C-PGC
Perturbation Generator Conditioned on Cross-modal knowledge. Previous generative universal attacks [17, 3] have shown excellent efficacy in fooling the discriminative models. Nevertheless, since existing generative attacks are limited to a single modality, directly utilizing the off-the-shelf generators might fail to leverage the multimodal interactions in these special V+L scenarios. To address this limitation, we additionally introduce cross-modal embeddings as auxiliary information to further facilitate the process of perturbation generation. Specifically, we modify the architecture of existing decoder-based generators by adding several cross-attention modules that have been proven effective in tasks with multiple input modalities. The obtained textual embeddings encoded by are then incorporated into our generator through:
@@
where is the flattened intermediate features within , and , , are the optimized parameters of attention modules.
@@
Multimodal Contrastive Loss. The preceding analysis regarding the failures of existing UAP attacks encourages us to design a loss function that can guide the generated UAP to break the learned multimodal feature alignment. Motivated by the fact that contrastive learning underpins the cross-modal alignment, we advocate leveraging this mechanism to attack VLP models themselves by contrastively training our C-PGC to essentially disrupt the benign alignment relationship. Concretely, we adopt the widely recognized InfoNCE [19] as our basic contrastive loss.
@@
To establish the contrastive paradigm, we first define the adversarial image as the anchor sample. Besides, it is also necessary to construct an appropriate set of positive and negative samples. Based on the fundamental objective of our attack, it is natural to leverage the originally matched text description set as negative samples to amplify the discrepancy between matched image-text pairs in the feature space of VLP models. Additionally, we need to select a set of positive samples to further pull the adversarial image away from its corresponding text descriptions . To this end, we propose a farthest selection strategy, which associates the anchor image with target texts whose embeddings differ significantly from that of the original clean image , to reach a more strong disruption of the multimodal alignment. Specifically, we randomly sample a batch of text sets from and select the text set with the largest feature distances from the current image as positive samples, i.e., . Moreover, we utilize data augmentations that resize the clean into diverse scales and apply random Gaussian noise to acquire a more diverse image set for set-level guidance [29]. With these well-constructed positive and negative samples, the multimodal contrastive loss can be formulated as:
@@
where is the universal image perturbation; , where denotes the temperature parameter and represents the cosine similarity.
@@
Unimodal Distance Loss. Apart from the multimodal guidance, we also consider the unimodal influence by directly pushing adversarial images away from their initial visual semantic area to improve the attack. Similarly, the input image also undergoes resizing and noise perturbation to generate the augmented image set for set-level guidance. Then, we craft the adversarial image through and process with the same augmentation operation to obtain the adversarial image set . Finally, we minimize the negative Euclidean distance between the embeddings of adversarial images and clean images to optimize the UAP generator. Formally, the loss is formulated as:
@@
Taking advantage of the unimodal guidance, ensures an effective optimization direction during the generator training and further enhances the attack effectiveness.
@@
Training Objective. With the above two well-designed loss terms and , the overall optimization objective of our conditional generator for image attacks can be formulated as:
@@
where is the pre-defined hyperparameter to balance the contributions of and . By training the network with the proposed loss function over the entire data distribution of the multimodal training dataset , the generator is optimized to generate UAPs that push the features of mismatched image-text pairs together while pulling the embeddings of the matched ones apart. This finally enables the generation of UAPs with strong generalization capabilities and high adversarial transferability.
@@
Text Modality Attacks. In textual attacks, the generator architecture and training loss are completely symmetrical with those of image attacks. Correspondingly, embeddings of the matched image are used as the cross-modal conditions for the generator. Given an adversarial text as the anchor sample, we use the set scaled from the originally matched image as negative samples while the augmented from the farthest image within the randomly sampled image set as positive samples to formulate the loss. is consequently calculated as the negative Euclidean distance between the embeddings of and the clean input . Accordingly, the conditional generator is utilized to output the adversarial textual embeddings, which are subsequently mapped back to the vocabulary space to obtain a universally applicable word-level perturbation.